For the Damast Corp. in 2014:
Income Statement for (in $) : 2014
Net sales : 372,300
Cost of goods sold : 156,000
Gross margin : 186,000
All sales are on credit.

Days of Accounts Receivable : 36.00
Accounts Receivable on Dec 31, 2013 : 46,000
How much were the Accounts Receivable on Dec 31, 2014?Calibri

23,324

24,696

26,068

27,440

28,812
Answer #1

Days of accounts receivable = 365 days * Average Account receivable / Net sales

36 days = 365 days * Average Account receivable / 372300

Average Account receivable = 36 * 372300 / 365

= 36720

Average account receivable = (Beginning receivable + Ending receivable) / 2

36720 = (46000 + Ending receivables) /2

36720 * 2 = 46000 + Ending receivables

Ending receivables = 73440 - 46000

= 27440

Correct choice D